# MultiCamCows2024 - A Multi-view Image Dataset for AI-driven Holstein-Friesian Cattle Re-Identification on a Working Farm

### v1

@@ 5 Discussions
Performance Improvements via Camera Integration. In addition, the self-supervised framework demonstrates superior performance using information from multiple cameras, outperforming their individual camera counterparts due to the enriched feature diversity for each label. Performance is improved compared to previous studies in similar settings [22, 23]. Conversely, the collective performance of the supervised classifier over the amalgamated camera dataset, while improved compared to some individual camera views, does not consistently surpass the performance when data from all cameras are used. As mentioned previously, the successful self-supervised training of the framework eliminates, the most time-consuming part of human labelling – that is labelling individuals – altogether. This, together with the performance gains made by merging data from disparate cameras, improves the viability of the visual ID approach on a commercial farm.
@@ 6 Conclusions and Future Work
Open-Set Implementation. In target re-identification applications, one of the main issues is that of how to deal with new individuals that were never part of the original training set. This happens on dairy farms fairly frequently as heifers enter the milking herd. Identifying possible new individuals is known as the open-set problem [57, 25, 58, 59], and solving this is an open challenge. In such scenarios, manual re-labelling of these unknown entities is not desirable. One possible solution to this problem is to build a classifier capable of deciding whether a cow is indeed ‘new’ or not, as in [60].
@@
7 Extras
Acknowledgements We acknowledge the support provided by Axel Montout in data availability and collection, as well as all other staff from the John Oldacre Centre at the Bristol Vet School to carry out this research project. For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) licence to any Author Accepted Manuscript version arising from this submission.
Funding This research was funded by the Biotechnology and Biological Sciences Research Council grant number BB/X017559/1. The APC was funded by UK Research and Innovation.
Institutional Review The animal study protocol was approved by the University of Bristol Animal Welfare Ethical Review Board under the code UIN-23-060 on the 37th September 2023.
Data Availablility The image dataset and raw camera footage is available at https://phoenix4582.github.io/MultiCamCows2024.github.io/
Conflict of Interest The authors declare no conflicts of interest.

### v2

@@ 5 Discussions
Performance Improvements via Camera Integration. In addition, the self-supervised framework demonstrates superior performance using information from multiple cameras, outperforming their individual camera counterparts due to the enriched feature diversity for each label. Performance is improved compared to previous studies in similar settings [14, 15]. From the baseline of our study, the same methodology can be potentially expanded towards camera systems of diverse heights, angles and perspectives. Conversely, the collective performance of the supervised classifier over the amalgamated camera dataset, while improved compared to some individual camera views, does not consistently surpass the performance when data from all cameras are used. As mentioned previously, the successful self-supervised training of the framework eliminates, the most time-consuming part of human labelling – that is labelling individuals – altogether. This, together with the performance gains made by merging data from disparate cameras, improves the viability of the visual ID approach on a commercial farm. Additionally, despite contributing a significantly larger number of images and introducing dataset imbalance, Camera 3 exhibited noticeably poorer performance compared to the combined views. This analysis demonstrates that quantity imbalance alone does not necessarily improve single-camera performance over combined views, as the diversity of image features plays a crucial role in identification accuracy.
[…]
@@ 6 Conclusions and Future Work
Open-Set Implementation. In target re-identification applications, one of the main issues is that of how to deal with new individuals that were never part of the original training set. This happens on dairy farms fairly frequently as heifers enter the milking herd. Identifying possible new individuals is known as the open-set problem [27, 28, 25], and solving this is an open challenge. In such scenarios, manual re-labelling of these unknown entities is not desirable. One possible solution to this problem is to build a classifier capable of deciding whether a cow is indeed ‘new’ or not, as in [29].